OpenStudy (anonymous):

What is 4 2/7*2 1/5? A. 8 2/35 B. 8 1/7 C. 9 1/7 D. 9 3/7

OpenStudy (bibby):

is that \[4 \frac{ 2 }{ 7 }\] and 2 \[\frac{ 1 }{ 5 }\]?

OpenStudy (anonymous):

yes

OpenStudy (bibby):

the first thing to do is to express the mixed fractions as improper fractions Do you remember how? Hint\[2 \frac{ 1 }{ 3 } = \frac{ 6+1 }{ 3 } = \frac{ 7 }{ 3 }\]:

OpenStudy (anonymous):

I think.

OpenStudy (bibby):

Here I'll write one of them out: \[4 \frac{ 2 }{ 7 } = \frac{ 7*4 }{ 7 } + \frac{ 2 }{ 7 } = \frac{ 28+2 }{ 7 }\] Do you see what I'm doing? I'll explain anything you don't understand

OpenStudy (bibby):

To get a mixed fraction into improper form, you multiply the whole number (outside the fraction) with the denominator and add it to the original numerator. So first I multiplied 4 * 7. Then added 2. And put that number over the original 7

OpenStudy (anonymous):

Oh ya I see that. Your multiplying the denominator and the whole number and the answer to the numerator.

OpenStudy (bibby):

Do you want to try converting 2 1/5 to an improper fraction?

OpenStudy (anonymous):

\[2 \frac{1}{ 5 }= \frac{ 5*2}{ 5 }+\frac{ 1 }{ 5 }= \frac{ 10+1 }{ 5 }\]

OpenStudy (bibby):

yup yup yup and what is \[\frac{ 30 }{ 7 } * \frac{ 11 }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 330 }{ 35 }\]

OpenStudy (bibby):

a+, 100, whoop whoop That's correct. Just remember that you can't multiply out mixed fractions like that. You have to get 'em into this form

OpenStudy (anonymous):

YAY